A 734 kg car is stuck on a road, 990 kg is traveling at 12 m/s to the east collides with the rear side of the car which is stuck on the road. In your role as an expert witness, you inspect the scene and determine that, after the collision, the 734 kg car slides on the road at 5 m/s in a direction 30 degrees north of east. Ignoring the effects of friction during the collision, calculate the magnitude of the velocity (in m/s) of the 990 kg car after the collision


1
Expert's answer
2021-11-29T19:08:43-0500
"990(12)=990v_x+734(5)\cos{30}\\v_x=8.8\frac{m}{s}\\
0=990v_y-734(5)\sin{30}\\v_y=1.85\frac{m}{s}\\v=\sqrt{1.85^2+8.8^2}=9.0\frac{m}{s}"